# Confidential Prompting: Protecting User Prompts from Cloud LLM Providers

### v2

@@ Abstract
Our work tackles the challenge of securing user inputs in cloud-hosted large language model (LLM) serving while ensuring output invariance, model confidentiality, and compute efficiency. We introduce secure multi-party decoding (SMD), which leverages confidential computing to confine user prompts to a trusted execution environment (TEE), namely a confidential virtual machine (CVM), while allowing service providers to generate tokens efficiently. We also introduce a novel cryptographic method, prompt obfuscation (PO), to ensure robustness against reconstruction attacks on SMD. We demonstrate that our approach preserves both prompt confidentiality and LLM serving efficiency. Our solution can enable privacy-preserving cloud LLM serving that handles sensitive prompts, such as clinical records, financial data, and personal information.
@@ 1 Introduction
This paper aims at confidential prompting in cloud-hosted LLM services, which achieves the confidentiality of user prompts while meeting three crucial constraints for commercial LLM service: output invariance, model confidentiality, and compute efficiency. Output invariance guarantees that the output of the LLM service remains the same regardless of whether confidentiality measures are applied to the prompts, ensuring the utility and reliability of the service. Model confidentiality ensures that the LLM’s model weights are kept secret from users. Compute efficiency requires that the introduction of confidentiality measures does not significantly increase the service cost. As a first step, we consider the Honest-but-Curious (HBC) threat model, in which the LLM provider faithfully follows all computation steps for LLM inference but tries to extract as much information as possible from the user.
@@
As we will present in §2.2 and §2.4, none of existing solutions meet all of these constraints. For example, differentially private in-context learning (DP-ICL) [5, 6] and data anonymization [7], alter the output, thus violating output invariance. Techniques like edge inference [8] and multiparty computation (MPC) [9] protect input by processing them locally in part, but this approach requires sharing some LLM weights with the user, which undermines model confidentiality. Although fully homomorphic encryption (FHE) [10, 11] preserves both output invariance and model confidentiality, its extensive computational demands restrict its feasibility for LLMs. More recently, Confidential Computing (CC) [12] has emerged as a promising approach to secure cloud-based DNN [13, 14, 15], particularly through the use of confidential virtual machines (CVMs) with GPU support. Although CVMs can ensure output invariance and model confidentiality, their naive applications demand complete trust in the LLM provider or prove impractical for LLM serving, due to the large memory footprint of LLMs within user CVMs, which limits scalability due to GPU memory constraints, and to the lack of batch parallelism.
@@
Figure 1: Overview of Confidential Prompting by OSMD. A user and their CVM collaborate to protect the user’s prompt from the LLM provider by applying Secure Multi-party Decoding (SMD) and Prompt Obfuscation (PO). The user prepares their prompt on a local machine, while the user’s CVM and the LLM provider are hosted in the cloud. PO occurs within the CVM during the prefill phase, while SMD involves both the CVM and the LLM provider during decoding.
@@
This work presents a new approach to confidential prompting that uses CVMs in an efficient and scalable manner, illustrated by Figure 1. Our key insight is that the delivery of LLM involves two distinct phases: prefill and decode, and the latter dominates computational use. Our approach, called Obfuscated Secure Multi-party Decoding (OSMD), implements prefill inside per-user CVMs and keeps the resulting KV attention states (called private KV cache) there, because prefill requires the plaintext prompt and the private KV cache can be used to reconstruct the prompt. OSMD then implements decode mostly outside the CVMs, without disclosing the private KV cache, using a technique called Secure Multi-party Decoding (SMD) (§4).
SMD formulates token generation (decode) as a secure two-party computation with one party being the CVM and the other the LLM provider: only the latter needs the LLM weights for decoding. The CVM already has the private KV cache from prefill and no longer requires any LLM weights for decode. On the other hand, the LLM provider computes the KV attention states for the newly generated tokens as the public KV cache. During the interaction between the CVM and the cloud LLM, the CVM computes the private attention score ($A_{pvt}$) with the private KV cache, and provides it to the LLM provider without exposing the private states. This enables the LLM provider to compute the final attention score and generate new tokens, without compromising output invariance. This partitioning approach is efficient because (1) the CVM does not have to retain the LLM model weights after prefill, and (2) the LLM provider can batch process the public attention states in parallel.
OSMD overcomes two challenges to SMD. First, SMD is vulnerable to (semi) black-box reconstruction attacks that approximate the private states with observations of the interaction and compromise prompt confidentiality. To address this, we introduce a novel cryptographic method, prompt obfuscation (PO)5), inspired by chaffing and winnowing2.3). The user’s CVM generates fake n-grams that appear as authentic as the original segments in the user’s prompt. The CVM uses these fake n-grams to create a set of virtual prompts. The LLM provider is responsible for generating tokens for the authentic prompt and each virtual prompt in parallel, but the CVM only relays the generated tokens of the only authentic prompt back to the user. We offer a formal analysis of the security provided by PO in §5.3, as well as a memory-efficient algorithm to serve multiple virtual prompts within the CVM. Second, prefill, performed within the CVM, requires LLM weights and as a result, may jeopardize model confidentiality. OSMD addresses this issue by allowing the LLM provider to check the integrity of the transferred output tokens, ensuring that the malicious client cannot exfiltrate LLM weights by sending arbitrary bytes. The cloud vendor can also limit the outbound traffic to only what is necessary to transmit the generated tokens. See §4.1.
@@
In §6, we report an implementation of OSMD (SMD+PO) using PyTorch. In §7 we evaluate our prototype on NVIDIA H100 GPU, with GPU CC enabled, against two naive CVM-based solutions (2(b) and 2(c)). We show that OSMD achieves 5$\times$ better latency than the naive CVM-based solution that defends against cloud LLM providers and scales well to (1) the number of concurrent users, (2) the size of the model, and (3) the number of input/output tokens. In §8, we conclude our work, discuss its limitations and what it takes to achieve stronger protection. For instance, SMD requires a weaker adversary model, i.e., HBC, compared to serving entirely in CVMs. Also, the effectiveness of PO is conditional, although we can detect potential information leaks and alert the user of such cases. We believe that cloud LLM serving that is both privacy-preserving and efficient is important and timely. Our work marks the first step towards utilizing confidential computing for secure LLM serving, and we hope it will spark further discussion on confidential prompting and methods offering stronger protection. We open-source OSMD in github.com/yale-sys/confidential-prompting.
@@ 2 Background and Related Work
(d) Proposed approach (SMD)
Figure 2: Various LLM serving approaches. (a) LLM serving without confidential computing (CC) is the most efficient, but user prompts are exposed to both the cloud infrastructure and the LLM provider, as well as any adversaries that may compromise the cloud. (b) Serving multiple users with a single CVM can protect against malicious cloud provider and adversaries on the cloud, but prompts remain exposed to the LLM provider. (c) Assigning a separate CVM per user offers higher level of protection, from both cloud infrastructure and LLM provider, but it is inefficient since each CVM contains its own copy of the LLM. (d) SMD strikes a balance between security and efficiency by isolating only the user’s prompt (and private KV cache) within the CVM, while allowing batch processing across users.
We design Confidential Prompting on top of the Key-Value (KV) Cache mechanism, i.e., reusing the key-value attention states in autoregressive decoding. It leverages confidential computing techniques to protect the private KV cache. Its design of prompt obfuscation is inspired by the design of chaffing and winnowing. We next provide a succinct background about these mechanisms and techniques.
@@ 2.1 LLM Inference with KV cache
2.1.1 Tokens
A token is an atomic unit of text that the LLM processes, representing parts of words, entire words, or even punctuation. The exact breakdown into tokens depends on the model’s tokenization algorithm, which converts a string into a sequence of tokens. For example, a sentence like “The cat sat” might be tokenized as (“The”, “cat”, “sat”), while more complex words can be split into multiple tokens (e.g., “unbelievable” becomes (“un”, “believ”, “able”)). The typical size of a token vocabulary ranges from 12K to 128K.
@@ 2.1.2 LLM Inference
We consider GPT-style LLMs [16, 17, 18, 19], which are trained to predict the distribution of the next token, $s_{n+1}$, given a sequence of tokens $s_{1},\dots,s_{n}$, known as causal language modeling. This prediction process uses the Transformer architecture [20], which consists of multiple self-attention layers. For a sequence of length $n$, represented as $X\in\mathbb{R}^{n\times d}$, the Transformer produces an output sequence $Y\in\mathbb{R}^{n\times d}$, where $d$ is the hidden dimension size. The self-attention mechanism, a core component of the Transformer, requires five matrix multiplications. First, the model computes matrices $Q=XW_{Q}$, $K=XW_{K}$, and $V=XW_{V}$, where $W_{Q}$, $W_{K}$, and $W_{V}\in\mathbb{R}^{d\times d}$ are trainable weight matrices. Next, the output is calculated as $Y=\sigma(QK^{\top})V\in\mathbb{R}^{n\times d}$, where $\sigma(\cdot)$ denotes the softmax function. The output becomes an input to the next self-attention layer. The LLM samples the next token $s_{n+1}$ from the distribution and appends it to the token sequence, iteratively until some termination condition is met, so called autoregressive token generation.
@@ 2.1.3 KV Cache
The KV cache mechanism [21, 22, 23, 24] is a technique frequently used to improve LLM inference efficiency. This mechanism leverages the causal nature of LLMs: when predicting a token $s_{i}$ within an input token sequence, the attention computation only considers preceding tokens, $s_{1},\dots,s_{i-1}$, rather than any tokens that follow. Consequently, instead of recalculating attention for all tokens in the input at each generation step, the LLM inference engine “caches” previously computed attention states and reuses them for subsequent inferences. Since the reusable attention states consist of the $K$ and $V$ tensors for each token, this technique is referred to as KV caching.
@@ 2.2.1 Confidential Computing
Confidential computing protects data in use, complementing traditional security measures such as encryption that protect data at rest and data in transit. The most common approach to confidential computing is the use of trusted execution environments (TEEs), such as enclaves and confidential virtual machines (CVMs), which are provided by hardware features like AMD SEV [25] and Intel TDX [26]. These environments isolate sensitive data and code. Thanks to their strong isolation capabilities, hardware-based TEEs ensure that even privileged entities such as the OS and hypervisor cannot access the data being processed. In addition to isolation, most hardware-based solutions also provide runtime encryption and remote attestation. Runtime encryption guarantees that all data and code within a CVM (or enclave) are encrypted in memory, offering an additional layer of protection. Remote attestation, on the other hand, allows users to verify the integrity of a remote CVM (or enclave) running in the cloud before transmitting any sensitive data. Originally exclusive to CPUs, confidential computing is now supported by the latest GPUs as well, such as the NVIDIA H100 [27]. Many cloud providers offer confidential computing as a service, primarily through the use of CVMs [14]. Unlike regular VMs, which can share memory for read-only data such as LLM weights, CVMs do not share memory, to ensure that confidential data remains securely isolated and encrypted. In this work, we use the terms TEE and CVM interchangeably.
@@ 2.2.2 Confidential Inferencing
A straightforward approach to applying confidential computing to the LLM service is to serve multiple users with a single LLM instance running within a shared CVM (2(b)). This method is used by most existing commercial services, such as confidential inferencing in Azure [14] and AWS [15]. Although this approach effectively protects against potential threats from malicious cloud providers and adversaries that can compromise the cloud, it requires all users to fully trust the LLM provider and the CVM owner, which is not aligned with our threat model (§3.2). An alternative approach is to instantiate a separate LLM instance within a CVM for each user (2(c)). This setup offers stronger security, as each user is the owner of their associated CVM, making it preferable in scenarios with only a few users, as discussed in §8. However, this approach suffers from two significant inefficiencies: (1) limited scalability, as the number of concurrent user CVMs is constrained, and (2) low throughput and high latency due to reduced batch parallelism. For example, a 13B-parameter LLM requires about 27 GB of memory for its weights using 16-bit floating point, which means that a 80 GB GPU can only support up to three concurrent user CVMs. Moreover, inference is performed independently for each user, e.g., $X_{1}W,\cdots,X_{n}W$, which is less efficient than batching as $(X_{1}:\cdots:X_{n})W$. Our solution, SMD, addresses both challenges by isolating only the user prompt within the CVMs while sharing the LLM among multiple users (2(d)).
2.3 Confidentiality without Encryption
In an information-theoretic sense, a message is secure if an eavesdropper can do no better than guessing randomly to determine its content [28, 29]. Modern encryption achieves this by transforming messages into ciphertexts that are indistinguishable from random noise. However, encryption is not the only method for ensuring message confidentiality. In this section, we first introduce the concept of chaffing and winnowing [30, 31] as an alternative approach to secure message confidentiality, and then present our security model for formal analysis in the following sections.
2.3.1 Chaffing and Winnowing

### v3

@@ Abstract
Our work tackles the challenge of securing user inputs in cloud-hosted large language model (LLM) serving while ensuring model confidentiality, output invariance, and compute efficiency. We introduce Secure Partitioned Decoding (SPD), which uses confidential computing to confine user prompts to a trusted execution environment (TEE), namely a confidential virtual machine (CVM), while allowing service providers to generate tokens efficiently. We also introduce a novel cryptographic method, Prompt Obfuscation (PO), to ensure robustness against reconstruction attacks on SPD. We demonstrate our approach preserves both prompt confidentiality and LLM serving efficiency. Our solution enables privacy-preserving cloud LLM serving that handles sensitive prompts, such as clinical records, financial data, and personal information.
@@ 1 Introduction
This paper aims at confidential prompting in cloud-hosted LLM services, which secures the confidentiality of user prompts while achieving another three crucial goals for commercial LLM service: model confidentiality, output invariance, and compute efficiency. (i) Model confidentiality ensures the LLM weights are kept secret from users. (ii) Output invariance guarantees the output of the LLM service remains the same regardless of whether confidentiality measures are applied, ensuring the utility and reliability of the service. (iii) Compute efficiency requires the introduction of confidentiality measures does not significantly increase the service cost.
We assume the users and the cloud LLM providers are mutually untrusted and potentially seek to uncover each other’s secrets, namely user prompts and LLM weights. In particular for the cloud LLM providers, we consider two variants of Honest-but-Curious (HBC) threat model, which are HBC without or with unlimited computational resources. In both scenarios, the service providers adhere strictly to the prescribed computation steps for LLM inference; however, they also endeavor to extract user secrets with various resource budgets. Our approach provides different security guarantees over user prompt confidentiality under different HBC variants, specifically securing the entire prompts or sensitive segments in the prompts such as Personal Identifiable Information (PII).
@@
As we present in §2.2 and §2.4, none of existing solutions achieve all of our goals under the same threat model. For example, techniques like edge inference [30] protect input by processing them locally, which do not work for large models hosted on cloud and require sharing LLM weights with the users, undermining model confidentiality. Differentially private in-context learning (DP-ICL) [59, 54] and data anonymization [52, 60, 10, 24], alter the output, thus violating output invariance. Although fully homomorphic encryption (FHE) [22, 19] preserves both model confidentiality and output invariance, its extensive computational demands restrict its feasibility for LLMs. More recently, Confidential Computing (CC) [8] has emerged as a promising approach to secure cloud-based DNN [26, 49, 47], particularly through the use of confidential virtual machines (CVMs) with GPU support. Although CVMs ensure output invariance, their naive applications either demand complete trust in the LLM provider or prove impractical for LLM serving. This impracticality raises due to the lack of batch parallelism and the large memory footprint of LLMs in user CVMs, which limits the scalability.
@@
Figure 1: Overview of Confidential Prompting by OSPD. A user and its CVM collaborate to protect the user’s prompt from the cloud LLM provider by applying Secure Partitioned Decoding (SPD) and Prompt Obfuscation (PO). The user sends its encrypted prompt to its CVM hosted in the cloud. Then virtual prompts and private KV cache are prepared within the CVM during the prefill phase, while SPD involves both the CVM and the cloud LLM provider during the decode phase.
@@
This work presents a new approach to confidential prompting that uses CVMs in an efficient and scalable manner, illustrated by Figure 1. Our key insight is that the delivery of LLM involves two distinct phases: prefill and decode, and the latter dominates computational use. Our approach, called Obfuscated Secure Partitioned Decoding (OSPD), performs prefill inside per-user CVMs and keeps the resulting KV attention states (called private KV cache) there, as prefill requires the plaintext prompt and the private KV cache can be used to extract the prompt. OSPD then performs decode mostly outside the CVMs, without disclosing the private KV cache, using a technique called Secure Partitioned Decoding (SPD) (§4).
SPD formulates token generation (decode) as a secure two-party computation with one party being the CVM and the other the cloud LLM provider. In other words, we partition the attention score computation into two parts: the private and the public attention scores ($A_{pvt}$ and $A_{pub}$), which are computed by the two parties respectively. During the decode, the CVM uses the private KV cache that is already precomputed in prefill to compute $A_{pvt}$. It does not require the LLM weights for computing $A_{pvt}$, thereby reducing the memory footprint of the CVM in decode. On the other hand, the cloud LLM providers compute the $A_{pub}$, merges it with $A_{pvt}$ received from the CVM to generate new tokens, and then maintains the public KV cache for these newly generated tokens.
SPD secures the entire user prompt when assuming the attackers are HBC without unlimited computational resources. The user prompt is secure as the CVM exposes only the $A_{pvt}$, which does not expose the prompts or the private KV cache. SPD is also efficient because the service providers can batch process the public KV cache for multiple users in parallel, and the CVM does not retain LLM weights in decode, thereby remaining small footprint. Finally, SPD ensures output invariance as the attention score decomposition is lossless (§4.2).
However, SPD leaves three unsolved challenges. First, SPD is vulnerable to (semi) black-box reconstruction attacks, which approximate the prompt with public information to get user secrets if an attacker has unlimited computational resources. Second, SPD does not protect the LLM responses, unlike the naive CVM approaches that process completely within the CVM (§2.2.2). It is worth noting that, attempts to uncover a prompt with its response is actually a kind of reconstruction attack. Details are in §4.3. Lastly, prefill requires LLM weights within the CVM and as a result, may jeopardize model confidentiality. OSPD mitigates the first two challenges with Prompt Obfuscation (PO) (§5) and solve the last challenge with control over the cloud infrastructure.
PO is a novel cryptographic method inspired by chaffing and winnowing2.3). To protect sensitive segments in a prompt, such as Personal Identifiable Information (PII), the CVM generates fake n-grams that appear as authentic as the original segments in the prompt. The CVM uses these fake n-grams to create $\lambda$ virtual prompts and then process the $\lambda+1$ prompts with prefill and SPD in parallel. In other words, the service providers learn nothing about the authentic and the virtual prompts or their private KV cache, except receiving ($\lambda$+1) private attention scores in a batch. Only the user and its CVM know which prompt (and response) is authentic.
PO mitigates the first two unsolved challenges left by SPD instead of offering absolute security. PO obfuscates the cloud LLM providers, but the latter may still succeed in guessing the authentic prompt (and response) with probability of $\epsilon+1/(\lambda+1)$, where $\epsilon$ is an error bound parameter used in sampling fake n-grams (§5.1). We offer a security analysis of PO in §5.2.
SPD and PO are complementary in two ways. On one hand, PO obfuscates the providers and offers a certain level of protection against reconstruction attacks. This minimizes the risk of leaking sensitive information in user prompts, while allowing SPD to offload majority of computation to out of CVMs. In other words, PO allows SPD to trade efficiency and scalability with minimal risk compared to naive CVM approaches. On the other hand, like many existing obfuscation methods, a knowledgeable attacker with more statistical information over the prompts has higher probability to break PO’s protection. However, SPD adds computational cost for the attacker to collect such statistical information and reduces its confidence in the collected data. Even if the attacker performs reconstruction attacks with unlimited computational resources, it hardly learns the exact plaintext prompts.
As for model confidentiality, OSPD addresses it by allowing the cloud provider to check the integrity of the transferred output tokens, ensuring that the malicious client cannot exfiltrate LLM weights by sending arbitrary bytes. The cloud provider can also limit the outbound traffic to only what is necessary to transmit the generated tokens (§4.1).
@@
In §6, we report an implementation of OSPD (SPD+PO) using PyTorch. In §7 we evaluate our prototype on NVIDIA H100 GPU with CC enabled, compared with two naive CVM approaches (§2.2.2). We show that OSPD scales well to the number of concurrent users and achieves 5$\times$ better latency than the naive CVM approach that defends against cloud LLM providers. We conclude our work in §8 and discuss its limitations in §A.4, such as how to achieve stronger protection. We believe that cloud LLM serving that is both privacy-preserving and efficient is important and timely. Our work marks the first step towards utilizing confidential computing for secure LLM serving, and we hope it will spark further discussion on confidential prompting and methods offering stronger protection. The source code for OSPD can be found at github.com/yale-sys/confidential-prompting.
@@ 2 Background and Related Work
(d) Proposed approach (SPD)
Figure 2: Various LLM serving approaches. (a) LLM serving without confidential computing (CC) is the most efficient, but user prompts are exposed to both the cloud infrastructure and the LLM provider. (b) Serving multiple users with a single CVM can protect against malicious cloud provider and adversaries on the cloud, but prompts remain exposed to the LLM provider. (c) Assigning a dedicated CVM per user offers higher level of protection, from both cloud infrastructure and LLM provider, but it is inefficient since each CVM contains its own copy of the LLM. (d) SPD strikes a balance between security and efficiency by isolating only the user’s prompt (and private KV cache) within the CVM, while allowing batch processing across users.
We next provide a succinct background about related techniques.
@@ 2.1.1 LLM Inference
We consider GPT-style LLMs [45, 7, 1, 55], which are trained to predict the distribution of the next token, $x_{n+1}$, given a sequence of tokens $x_{1},\dots,x_{n}$, known as causal language modeling. This prediction process uses the Transformer architecture [56], which consists of multiple self-attention layers. For a sequence of length $n$, represented as $X\in\mathbb{R}^{n\times d}$, the Transformer produces an output sequence $Y\in\mathbb{R}^{n\times d}$, where $d$ is the hidden dimension size. The self-attention mechanism, a core component of the Transformer, requires five matrix multiplications. First, the model computes matrices $Q=XW_{Q}$, $K=XW_{K}$, and $V=XW_{V}$, where $W_{Q}$, $W_{K}$, and $W_{V}\in\mathbb{R}^{d\times d}$ are trainable weight matrices. Next, the output is calculated as $Y=\sigma(QK^{\top})V\in\mathbb{R}^{n\times d}$, where $\sigma(\cdot)$ denotes the softmax function. The output becomes an input to the next layer. The LLM samples the next token $x_{n+1}$ from the distribution and appends it to the token sequence, iteratively until some termination condition is met, so called autoregressive token generation.
@@ 2.1.2 KV Cache
The KV cache mechanism [41, 53, 44, 18] is a technique frequently used to improve LLM inference efficiency. This mechanism leverages the causal nature of LLMs: when predicting a token $x_{i}$ within a token sequence, the attention computation only considers preceding tokens, $x_{1},\dots,x_{i-1}$, rather than any tokens that follow. Consequently, instead of recalculating attention for all tokens at each generation step, the LLM inference engine “caches” previously computed attention states and reuses them for subsequent inferences. Since the reusable attention states consist of the $K$ and $V$ matrices for each token, this technique is referred to as KV caching.
@@ 2.2.1 Confidential Computing
Confidential computing protects data in use, complementing traditional security measures such as encryption that protect data at rest and data in transit. The most common approach to confidential computing is to use trusted execution environments (TEEs), such as enclaves and confidential virtual machines (CVMs), which are provided by hardware features like AMD SEV [3] and Intel TDX [23]. These environments isolate sensitive data and code. Thanks to their strong isolation capabilities, hardware-based TEEs ensure even privileged entities such as the OS and hypervisor cannot access the data being processed. In addition to isolation, most hardware-based TEEs also provide runtime encryption and remote attestation. Runtime encryption guarantees all code and data within a CVM (or enclave) are encrypted in memory, offering an additional layer of protection. Remote attestation allows users to verify the integrity of a remote CVM (or enclave) running in the cloud before transmitting any sensitive data. Originally exclusive to CPUs, confidential computing is now supported by the latest GPUs as well, such as the NVIDIA H100 [40]. Many cloud providers offer confidential computing as a service, primarily through the use of CVMs [49]. Unlike regular VMs, which can share memory for read-only data such as LLM weights, CVMs do not share memory, to ensure confidential data remains securely isolated and encrypted. In this work, we use the terms TEE and CVM interchangeably.
@@ 2.2.2 Confidential Inferencing
An intuitive approach to apply confidential computing to the LLM service is serving multiple users with a single LLM instance running in a shared CVM (2(b)). This method is used by most existing commercial services, such as confidential inferencing in Azure [49] and AWS [47]. Although this approach effectively protects against malicious cloud providers and adversaries that compromise the cloud, it requires all users to fully trust the LLM provider and the CVM owner, which is not aligned with our threat model (§3.2).
@@
An alternative approach is to instantiate a dedicated LLM instance for per-user CVM (2(c)). This setup offers stronger security, as each user is the owner of its associated CVM. However, this approach suffers from two significant inefficiencies: (1) limited scalability, as the number of concurrent user CVMs is constrained, and (2) low throughput due to reduced batch parallelism. For example, a 13B-parameter LLM requires about 27 GB of memory for its weights using 16-bit floating point, which means that a 80 GB GPU can only support up to three concurrent user CVMs. Moreover, inference is performed independently for each user, e.g., $X_{1}W,\cdots,X_{n}W$, which is less efficient than batching as $(X_{1}:\cdots:X_{n})W$. Our solution, SPD, addresses both challenges by isolating only the user prompt within the CVM while sharing the LLM among multiple users (2(d)). However, we note SPD alone is less secure as it offloads majority of computation to out of CVMs, trading efficiency and scalability with minimal risk of secret leakage. Its goal is not to replace confidential inferencing solutions but to provide an alternative approach that fits different scenarios as discussed in §A.4.
Some prior work [51, 61, 57] aim to address the scalability problem by isolating multiple instances within a single CVM. These approaches either fail to use GPUs for acceleration, or do not provide the same security level as all instances must trust and share the software stack deployed within the CVM.
2.3 Chaffing and Winnowing